1 kg of steam at 100 bar and 370℃ expands reversibly in a perfectly thermally insulated cylinder behind a piston until the pressure is 38 bar and the steam is then dry saturated. Calculate the work done


1
Expert's answer
2022-03-30T12:51:02-0400

"U=H-PV"

"U_1=2998.82-\frac{100*0.002416*10^5}{10^3}=2757.22 (KJ/kg)"

"U_g=2801.18-\frac{38*0.0524672*10^5}{10^3}=2601.8 (KJ/kg)"

"W=U_1-U_2=2757.22-2601.8=155.42(KJ\/kg)"